It is proposed to build refrigeration plant for a cold storage to be maintained at -3°C. The ambient temperature is 27°C. If 5×106 kJ/h of energy is to be continuously removed from the cold storage, the minimum power required to run the refrigerator will be
14.3Kw

Grade:12th pass

1 Answers

Susmita
5 years ago
I hope you know the relation
(Q1/T1)=(Q2/T2)
where Q1 =heat rejected in hot reservoir=5×106kJ
T=temperature of hot reservoir=(27+273)K=300K
 Q=heat absorbed from cold reservoir
T2=temperature of cold reservoir=(-3+273)=300K
Ussing the relation and putting the values
Q2 =5.55×10kJ
Work input=Q1-Q2=0.55 ×106kJ
Power=0.55kJ/h=(0.55×106)/(60×60) kJ/s=154.3kJ/s=154.3 kW